Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
4x + 8y = 16
-2x + 5y = 55

Answer by mananth(16949)

4 x + 8 y = 16 .............1
-2 x + 5 y = 55 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 8
-20 x -40 y = -80
-16 x 40 y = 440
Add the two equations
-36 x = 360
/ -36
x = -10
plug value of x in (1)
4 x + 8 y = 16
-40 + 8 y = 16
8 y = 16 + 40
8 y = 56
y = 7
Solution (-10,7)